# Improving Adversarial Robust Fairness via Anti-Bias Soft Label Distillation

### v1

@@ Abstract
Adversarial Training (AT) has been widely proved to be an effective method to improve the adversarial robustness against adversarial examples for Deep Neural Networks (DNNs). As a variant of AT, Adversarial Robustness Distillation (ARD) has demonstrated its superior performance in improving the robustness of small student models with the guidance of large teacher models. However, both AT and ARD encounter the robust fairness problem: these models exhibit strong robustness when facing part of classes (easy class), but weak robustness when facing others (hard class). In this paper, we give an in-depth analysis of the potential factors and argue that the smoothness degree of samples’ soft labels for different classes (i.e., hard class or easy class) will affect the robust fairness of DNN models from both empirical observation and theoretical analysis. Based on the above finding, we propose an Anti-Bias Soft Label Distillation (ABSLD) method to mitigate the adversarial robust fairness problem within the framework of Knowledge Distillation (KD). Specifically, ABSLD adaptively reduces the student’s error risk gap between different classes to achieve fairness by adjusting the class-wise smoothness degree of samples’ soft labels during the training process, and the smoothness degree of soft labels is controlled by assigning different temperatures in KD to different classes. Extensive experiments demonstrate that ABSLD outperforms state-of-the-art AT, ARD, and robust fairness methods in terms of overall performance of robustness and fairness.
@@ 1 Introduction
Figure 1: The comparison between the existing sample-based fair adversarial training and our label-based fair adversarial training. For the former ideology in (a), the trained model’s bias is avoided by re-weighting the sample’s importance according to their different contribution to class-wise fairness. For the latter ideology in (b), the trained model’s bias is avoided by re-temperating the smoothness degree of soft labels for different types of classes.
@@
Deep neural networks (DNNs) have achieved great success in various tasks, e.g., classification [10], detection [6], and segmentation [22]. However, DNNs are vulnerable to adversarial attacks [27, 33, 31, 32], where adding small perturbations to the input examples will lead to misclassification. To enhance the robustness of DNNs, Adversarial Training (AT) [18, 39, 30, 15] is proposed and has been proven to be an effective method to defend against adversarial examples. AT can be formulated as a min-max optimization, where the inner maximization can be regarded as generating adversarial examples, and outer minimization is to train the model by adversarial examples generated in maximization. To further improve the robustness of small DNNs, Adversarial Robustness Distillation (ARD) [7] as a variant of AT is proposed and aims to transfer the robustness of the large teacher models into the small student models based on Knowledge Distillation (KD), and further researches [42, 43, 40, 12, 41] demonstrate the excellent performance of ARD in improving the robustness of small models.
@@
Although AT and ARD can remarkably improve the adversarial robustness, some researches [2, 29, 37, 17, 26, 36] demonstrate the robust fairness problem: these models perform strong robustness on part of classes (easy class) but show high vulnerability on others (hard class). This phenomenon will raise further attention to class-wise security. Specifically, an overall robust model appears to be relatively safe for model users, however, the robust model with poor robust fairness will lead to attackers targeting vulnerable classes of the model, which leads to significant security risks to potential applications. Different from simply improving the overall robustness, some methods are proposed to address the robust fairness problem in AT and ARD [37, 34, 36, 17, 26] (i.e., increasing the worst class robustness as much as possible while keeping the overall robustness constant) and have achieved some progress. However, the robust fairness problem still exists and requires further to be explored.
@@
In this paper, we give an in-depth analysis of the potential factors to influence robust fairness in the optimization objective function. From the perspective of the training sample, the sample itself has a certain degree of biased behavior, which is mainly reflected in the different learning difficulties and various vulnerabilities to adversarial attacks. For this reason, previous works usually apply the re-weighting ideology to achieve robust fairness for different types of classes in the optimization process [37, 34, 36]. However, as another important factor in the optimization objective function, the role of the labels applied to guide the model’s training is ignored. Labels can be divided into two types, including one-hot labels and soft labels, where the soft labels are widely studied [28, 11] and have been proven effective in improving the performance of DNN models. Inspired by this, we try to explore robust fairness from the perspective of samples’ soft labels. Interestingly, we find that the smoothness degree of soft labels for different classes (i.e., hard class and easy class) can affect the class-wise robust fairness of DNN models from both empirical observation and theoretical analysis. Specifically, it is helpful to improve robust fairness by assigning sharp soft labels for hard classes and smooth soft labels for easy classes.
@@
Based on the above finding, we further propose an Anti-Bias Soft Label Distillation (ABSLD) method to mitigate the adversarial robust fairness problem within the framework of knowledge distillation. ABSLD can adaptively adjust the smoothness degree of soft labels by re-temperating the teacher’s soft labels for different classes, and each class has its own teacher’s temperatures based on the student’s error risk. For instance, when the student performs more error risk in some classes, ABSLD will compute sharp soft labels by assigning lower temperatures, and the student’s learning intensity for these classes will relatively increase compared with other classes. After the optimization, the student’s robust error risk gap between different classes will be reduced.
@@
  • We explore the labels’ effects on the adversarial robust fairness of DNN models, which is different from the existing sample perspective. To the best of our knowledge, we are the first one to find that the smoothness degree of samples’ soft labels for different types of classes can affect the robust fairness from both empirical observation and theoretical analysis.
@@
  • We empirically verify the effectiveness of ABSLD. Extensive experiments on different datasets and models demonstrate that our ABSLD can outperform state-of-the-art methods against a variety of attacks in terms of overall performance of robustness and fairness.
@@ 2.1 Adversarial Training
To defend against the adversarial examples, Adversarial Training (AT) [18, 39, 30, 13, 23] is regarded as an effective method to obtain robust models. AT can be formulated as a min-max optimization problem as follows:
@@
where $f(\cdot;\theta)$ represents a deep neural network with weight $\theta$, $D$ represents a data distribution with clean example $x$ and the ground truth label $y$. $\mathcal{L}$ represents the optimization loss function, e.g. the cross-entropy loss. $\delta$ represents the adversarial perturbation, and $\Omega$ represents a bound, which can be defined as $\Omega=\left\{\delta:||\delta||\leq\epsilon\right\}$ with the maximum perturbation scale $\epsilon$. To generate adversarial examples for the inner maximization in Eq.(1), Projected Gradient Descent (PGD) [18] is regarded as a common approach in AT, which takes multiple gradient steps as follows:
$\delta^{n+1}=\Pi_{\Omega}\big{(}\delta^{n}+\alpha\cdot sign(\nabla_{x}\mathcal{L}(f(x+\delta^{n};\theta),y))\big{)},$ (2)
where $\delta^{n}$ is the adversarial perturbation at the $n$-th step, $\Pi(\cdot)$ is the projection function, and $\alpha$ is a step size. $\delta^{0}$ is initialized as a uniform distribution limited in $\Omega$. For the PGD with total step num $N$, the final adversarial perturbation is given by $\delta^{*}=\delta^{N}$. To further improve the performance, some variant methods of AT appear including regularization [19, 39, 30], using additional data [25, 20], and optimizing iteration process [14, 21]. Different from the above methods for improving the overall robustness, in this paper, we mainly focus on solving the robust fairness problem.
@@ 2.2 Adversarial Robustness Distillation
Knowledge Distillation [11] as a training method can effectively transfer the large model’s knowledge into the large model’s knowledge, which has been widely applied in different areas. To enhance the adversarial robustness of small DNNs, Goldblum et al. [7] first propose Adversarial Robustness Distillation (ARD) by applying the clean prediction distribution of strong robust teacher models to guide the adversarial training of student models. Zhu et al. [42] argue that the prediction of the teacher model is not so reliable, and composite with unreliable teacher guidance and student introspection during the training process. RSLAD [43] applies the teacher clean prediction distribution as the guidance to train both clean examples and adversarial examples. MTARD [40, 41] applies clean teacher and adversarial teacher to enhance both accuracy and robustness, respectively. AdaAD [12] adaptively searches for optimal match points by directly applying the teacher adversarial prediction distribution in the inner maximization.
In this paper, we explore how to enhance robust fairness within the framework of knowledge distillation.
@@ 2.3 Adversarial Robust Fairness
Some researchers address the robust fairness problem of AT from different views [37, 34, 36, 17, 35, 26]. The most intuitive idea is to give different weights to the sample of different classes in the optimization process, and Xu et al. [37] propose Fair Robust Learning (FRL), which adjusts the loss weight and the adversarial margin based on the prediction accuracy of different classes. Ma et al. [17] propose Fairly Adversarial Training (FAT) to mitigate the trade-off between robustness and fairness by adding a regularization loss to control the variance of class-wise adversarial error risk. Sun et al. [26] propose Balance Adversarial Training (BAT) to achieve both source-class fairness (different difficulties in generating adversarial examples from each class) and target-class fairness (disparate target class tendencies when generating adversarial examples). Wu et al. [35] argue that the maximum entropy regularization for the model’s prediction distribution can help to achieve robust fairness. Wei et al. [34] propose Class-wise Calibrated Fair Adversarial Training (CFA) to address fairness by dynamically customizing adversarial configurations for different classes and modifying the weight averaging technique. To enhance the ARD robust fairness, Yue et al. [36] propose Fair-ARD by re-weighting different classes based on the vulnerable degree in the optimization process.
Different from these sample-based fairness methods, we try to solve this problem from the perspective of samples’ labels, by adjusting the class-wise smoothness degree of samples’ soft labels in the optimization process.
3 Robust Fairness versus Smoothness Degree of Soft Labels
As an important part of the model optimization process, label information used to guide the model plays an important role. The label can be divided into one-hot labels and soft labels, where one-hot labels only contain one class’s information and soft labels can be considered as an effective way to alleviate over-fitting and improve the performance [28]. Previous methods usually ignore the class-wise smoothness degree of soft labels, either applying the same smoothness degree [28], or uniformly changing the smoothness degree of soft labels for all the classes [11]. Different from previous methods, we are curious about if we artificially adjust the class-wise smoothness degree of soft labels, will it influence the class-wise robust fairness of the trained model? Here we try to explore the relationship between class-wise smoothness degree of soft labels and the robust fairness from both empirical observation and theoretical analysis.
@@ 3.1 Empirical Observation
Here, we mainly focus on the impact of the class-wise smoothness degree of soft labels on adversarial training. First, we train the model with soft labels that have the same smoothness degree for all types of classes. Then we assign different smoothness degrees of soft labels for hard classes and easy classes: specifically, we manually use sharper soft labels for hard classes and smoother soft labels for easy classes. In particular, we determine hard and easy classes based on prior knowledge. Here we conduct the experiment based on the SAT [18], and results are shown in Figure 2.
Figure 2: The class-wise and average robustness of DNN models guided by soft labels with the same smoothness degree (SSD) and different smoothness degree (DSD) for different classes, respectively. For the soft labels with different smoothness degrees, we use sharper soft labels for hard classes and use smoother soft labels for easy classes. We select two DNN models (ResNet-18 and MobileNet-v2) trained by SAT [18] on CIFAR-10. The robust accuracy is evaluated based on PGD. We see that blue lines and red lines have similar average robustness, but the worst robustness of blue lines are remarkably improved compared with red lines.
@@ 3.2 Theoretical Analysis
Here we try to theoretically analyze the impact of the smoothness degree of soft labels on class-wise fairness. Firstly, we want to analyze the model bias performance with the guidance of the soft label distribution with the same smoothness degree. Then we give Corollary 3.1 by extending the prediction distribution of binary linear classifier into the prediction distribution of DNNs based on the theoretical analysis in [37] and [17].
Corollary 3.1
@@
then the error risks for classes $c_{+}$ and $c_{-}$ have a relationship as follows:
@@
Corollary 3.1 actually demonstrates that when optimizing hard and easy classes with equal intensity, the model will inevitably be biased, and this bias mainly comes from the characteristics of the sample itself and is not related to the optimization method. Based on the Corollary 3.1, we can further analyze the performance differences with the guidance of the different types of soft labels. Here we provide Theorem 3.1 about the relationship between class-wise smoothness degree of soft labels and robust fairness.
Theorem 3.1
Follows the setting in Corollary 3.1, for a dataset $\mathcal{D}$ containing $2$ classes ($c_{+}$ and $c_{-}$), two soft label distribution ($P_{\lambda 1}=\{p_{c_{+}}^{\lambda 1},p_{c_{-}}^{\lambda 1}\}$ and $P_{\lambda 2}=\{p_{c_{+}}^{\lambda 2},p_{c_{-}}^{\lambda 2}\}$) exist, where $P_{\lambda 2}$ have a correct prediction distribution but have a limited different class-wise smoothness degree of soft labels ($v_{1}>0$, $v_{2}>0$):
@@
$p_{c_{+}}^{\lambda 1}(x_{c_{+}})>p_{c_{-}}^{\lambda 2}(x_{c_{-}})=p_{c_{+}}^{\lambda 1}(x_{c_{+}})-v_{2}>0.5,$ (7)
@@
$R(f(x_{c_{+}};$ $\theta_{\lambda 1}))-R(f(x_{c_{-}};\theta_{\lambda 1}))>$
$R(f(x_{c_{+}};\theta_{\lambda 2}))-R(f(x_{c_{-}};\theta_{\lambda 2})).$ (8)
The proof of Theorem 3.1 can be found in Appendix A. Based on Theorem 3.1, the class-wise smoothness degree of soft labels theoretically has an impact on class-wise robust fairness. If the soft label distribution with different smoothness degree $P_{\lambda 2}$ is applied to guide the model training, where the sharper smoothness degree of soft labels for hard classes and smoother smoothness degree of soft labels for easy classes, the model will appear smaller error risk gap between easy and hard class compared with the soft label distribution with same smoothness degree $P_{\lambda 1}$, which demonstrates better robust fairness. The Theorem 3.1 theoretically demonstrates that if we appropriately adjust the class-wise smoothness degree of soft labels, the model can achieve class-wise robust fairness.

### v2

@@ Abstract
Adversarial Training (AT) has been widely proved to be an effective method to improve the adversarial robustness against adversarial examples for Deep Neural Networks (DNNs). As a variant of AT, Adversarial Robustness Distillation (ARD) has demonstrated its superior performance in improving the robustness of small student models with the guidance of large teacher models. However, both AT and ARD encounter the robust fairness problem: these models exhibit strong robustness when facing part of classes (easy class), but weak robustness when facing others (hard class). In this paper, we give an in-depth analysis of the potential factors and argue that the smoothness degree of samples’ soft labels for different classes (i.e., hard class or easy class) will affect the robust fairness of DNNs from both empirical observation and theoretical analysis. Based on the above finding, we propose an Anti-Bias Soft Label Distillation (ABSLD) method to mitigate the adversarial robust fairness problem within the framework of Knowledge Distillation (KD). Specifically, ABSLD adaptively reduces the student’s error risk gap between different classes to achieve fairness by adjusting the class-wise smoothness degree of samples’ soft labels during the training process, and the smoothness degree of soft labels is controlled by assigning different temperatures in KD to different classes. Extensive experiments demonstrate that ABSLD outperforms state-of-the-art AT, ARD, and robust fairness methods in the comprehensive metric (Normalized Standard Deviation) of robustness and fairness.
@@ 1 Introduction
Figure 1: The comparison between the sample-based fair adversarial training and our label-based fair adversarial training. For the former ideology in (a), the trained model’s bias is avoided by re-weighting the sample’s importance according to the different contribution to fairness. For the latter ideology in (b), the trained model’s bias is avoided by re-temperating the smoothness degree of soft labels for different classes.
@@
Deep neural networks (DNNs) have achieved great success in various tasks, e.g., classification [10], detection [6], and segmentation [24]. However, DNNs are vulnerable to adversarial attacks [30; 35; 33; 34], where adding small perturbations to the input examples will lead to misclassification. To enhance the robustness of DNNs, Adversarial Training (AT) [20; 41; 32; 14] is proposed and has been proven to be an effective method to defend against adversarial examples. To further improve the robustness, Adversarial Robustness Distillation (ARD) [7] as a variant of AT is proposed and aims to transfer the robustness of the large models into the small models based on Knowledge Distillation (KD), and further researches [44; 45; 43; 12; 42] show the excellent performance of ARD.
@@
Although AT and ARD can remarkably improve the adversarial robustness, some researches [2; 31; 39; 19; 28; 38] demonstrate the robust fairness problem: these models perform strong robustness on part of classes (easy class) but show high vulnerability on others (hard class). This phenomenon will raise further attention to class-wise security. Specifically, an overall robust model appears to be relatively safe for model users, however, the robust model with poor robust fairness will lead to attackers targeting vulnerable classes of the model, which leads to significant security risks to potential applications. Different from simply improving the overall robustness, some methods are proposed to address the robust fairness problem in AT and ARD [39; 36; 38; 19; 28] (i.e., improving the worst-class robustness as much as possible without sacrificing too much overall robustness). However, the robust fairness problem still exists and requires further to be explored.
@@
For that, we give an in-depth analysis of the potential factors to influence robust fairness in the optimization objective function. From the perspective of the training sample, the sample itself has a certain degree of biased behavior, which is mainly reflected in the different learning difficulties and various vulnerabilities to adversarial attacks. For this reason, previous works apply the re-weighting ideology to achieve robust fairness for different types of classes in the optimization process [39; 36; 38]. However, as another important factor in the optimization objective function, the role of the labels applied to guide the model’s training is ignored. Labels can be divided into two types, including one-hot labels and soft labels, where the soft labels are widely studied [29; 11] and have been proven effective in improving the performance of DNNs. Inspired by this, we try to explore robust fairness from the perspective of samples’ soft labels. Interestingly, we first find that the smoothness degree of soft labels for different classes (i.e., hard and easy class) can affect the robust fairness of DNNs from both empirical observation and theoretical analysis. Intuitively speaking, sharper soft labels mean larger supervision intensity, while smoother soft labels mean smaller supervision intensity, so it is helpful to improve robust fairness by assigning sharp soft labels for hard classes and smooth soft labels for easy classes.
@@
Based on the above finding, we further propose an Anti-Bias Soft Label Distillation (ABSLD) method to mitigate the adversarial robust fairness problem within the framework of knowledge distillation. ABSLD can adaptively adjust the smoothness degree of soft labels by re-temperating the teacher’s soft labels for different classes, and each class has its own teacher’s temperatures based on the student’s error risk. For instance, when the student performs more error risk in some classes, ABSLD will compute sharp soft labels by assigning lower temperatures, and the student’s learning intensity for these classes will relatively increase compared with other classes. After the optimization, the student’s robust error risk gap between different classes will be reduced. The code can be found in https://github.com/zhaoshiji123/ABSLD.
@@
  • We explore the labels’ effects on the adversarial robust fairness of DNNs, which is different from the existing sample perspective. To the best of our knowledge, we are the first one to find that the smoothness degree of samples’ soft labels for different types of classes can affect the robust fairness from both empirical observation and theoretical analysis.
@@
  • We empirically verify the effectiveness of ABSLD. Extensive experiments on different datasets and models demonstrate that our ABSLD can outperform state-of-the-art methods against a variety of attacks in the comprehensive metric (Normalized Standard Deviation) of robustness and fairness.
@@ 2.1 Adversarial Training
To defend against the adversarial examples, Adversarial Training (AT) [20; 41; 32; 13; 25] is regarded as an effective method to obtain robust models. AT can be formulated as a min-max optimization problem as follows:
@@
where $f(\cdot;\theta)$ represents a deep neural network with weight $\theta$, $D$ represents a data distribution with clean example $x$ and the ground truth label $y$. $\mathcal{L}$ represents the optimization loss function, e.g. the cross-entropy loss. $\delta$ represents the adversarial perturbation, and $\Omega$ represents a bound, which can be defined as $\Omega=\left\{\delta:||\delta||\leq\epsilon\right\}$ with the maximum perturbation scale $\epsilon$. To further improve the performance, some variant methods of AT appear including regularization [21; 41; 32], using additional data [27; 22], and optimizing iteration process [15; 23]. Different from the above methods for improving the overall robustness, in this paper, we focus on solving the robust fairness problem.
@@ 2.2 Adversarial Robustness Distillation
Knowledge Distillation [11] as a training method can effectively transfer the large model’s knowledge into the small model’s knowledge, which has been widely applied in different areas. To enhance the adversarial robustness of small DNNs, Goldblum et al. [7] first propose Adversarial Robustness Distillation (ARD) by applying the clean prediction distribution of strong robust teacher models to guide the adversarial training of student models. Zhu et al. [44] argue that the prediction of the teacher model is not so reliable, and composite with unreliable teacher guidance and student introspection during the training process. RSLAD [45] applies the teacher clean prediction distribution as the guidance to train both clean examples and adversarial examples. MTARD [43; 42] applies clean teacher and adversarial teacher to enhance both accuracy and robustness, respectively. AdaAD [12] adaptively searches for optimal match points by directly applying the teacher adversarial prediction distribution in the inner maximization. In this paper, we explore how to enhance robust fairness within the framework of knowledge distillation.
@@ 2.3 Adversarial Robust Fairness
Some researchers address the robust fairness problem from different views [39; 18; 36; 38; 19; 37; 28] and improve the fairness without losing too much robustness. The most intuitive idea is to give different weights to the sample of different classes in the optimization process, and Xu et al. [39] propose Fair Robust Learning (FRL), which adjusts the loss weight and the adversarial margin based on the prediction accuracy of different classes. Ma et al. [19] finds the trade-off exists between robustness and fairness and propose Fairly Adversarial Training to mitigate this phenomenon by adding a regularization loss to control the variance of class-wise adversarial error risk. Sun et al. [28] propose Balance Adversarial Training (BAT) to achieve both source-class fairness (different difficulties in generating adversarial examples from each class) and target-class fairness (disparate target class tendencies when generating adversarial examples). Wu et al. [37] argue that the maximum entropy regularization for the model’s prediction distribution can help to achieve robust fairness. Wei et al. [36] propose Class-wise Calibrated Fair Adversarial Training (CFA) to address fairness by dynamically customizing adversarial configurations for different classes and modifying the weight averaging operation. To enhance the ARD robust fairness, Yue et al. [38] propose Fair-ARD by re-weighting different classes based on the vulnerable degree. Different from these sample-based fairness methods, we try to solve this problem from the perspective of samples’ labels, by adjusting the class-wise smoothness degree of samples’ soft labels in the optimization process.
3 Robust Fairness via Smoothness Degree of Soft Labels
As an important part of the model optimization, label information used to guide the model plays an important role. The label can be divided into one-hot labels and soft labels, where one-hot labels only contain one class’s information and soft labels can be considered as an effective way to alleviate over-fitting and improve the performance [29]. Previous methods usually ignore the class-wise smoothness degree of soft labels, either applying the same smoothness degree [29], or uniformly changing the smoothness degree of soft labels for all the classes without deliberate adjustments [11]. Different from previous methods, we are curious about if we adjust the class-wise smoothness degree of soft labels, will it influence the class-wise robust fairness of the trained model? Intuitively speaking, different smoothness degree of soft labels denote different supervision intensity, which means that it is possible to achieve fairness by adjusting the smoothness degree of soft labels. Here we try to explore the relationship between class-wise smoothness degree of soft labels and the robust fairness from both empirical observation and theoretical analysis.
@@ 3.1 Empirical Observation
Here, we focus on the impact of the class-wise smoothness degree of soft labels on adversarial training. First, we train the model with soft labels that have the same smoothness degree for all types of classes (smoothing coefficient [^1] For N-class one-hot ground truth labels, after processing by the smoothing coefficient $\gamma$, the highest probability (correct class) decreases from 1 to $1-\gamma$, and the other probability (wrong class) increases from 0 to $\frac{\gamma}{n-1}$. is 0.2). Then we assign different smoothness degrees of soft labels for hard classes and easy classes: specifically, we manually use sharper soft labels for hard classes (smoothing coefficient is 0.05) and smoother soft labels for easy classes (smoothing coefficient is 0.35). We conduct the experiment based on the SAT [20] shown in Figure 2.
(a) Class-wise and average Robustness of ResNet-18 on CIFAR-10.
(b) Class-wise and average Robustness of MobileNet-v2 on CIFAR-10.
Figure 2: The class-wise and average robustness of DNNs guided by soft labels with the same smoothness degree (SSD) and different smoothness degree (DSD) for different classes, respectively. For the soft labels with different smoothness degrees, we use sharper soft labels for hard classes and use smoother soft labels for easy classes. We select two DNNs (ResNet-18 and MobileNet-v2) trained by SAT [20] on CIFAR-10. The robust accuracy is evaluated based on PGD. The checkpoint is selected based on the best checkpoint of the highest mean value of all-class average robustness and the worst class robustness following [36]. We see that blue lines and red lines have similar average robustness, but the worst robustness of blue lines are remarkably improved compared with red lines.
@@ 3.2 Theoretical Analysis
Here we try to theoretically analyze the impact of the smoothness degree of soft labels on class-wise fairness. Firstly, we want to analyze the model bias performance with the guidance of the soft label distribution with the same smoothness degree. Then we give Corollary 1 by extending the prediction distribution of binary linear classifier into the prediction distribution of DNNs based on the theoretical analysis in [39] and [19].
Corollary 1.
@@
then the error risks (the expectation that samples are wrongly predicted by the model) for classes $c_{+}$ and $c_{-}$ have a relationship as follows:
@@
Corollary 1 demonstrates that when optimizing hard and easy classes with equal intensity, the model will inevitably be biased, and this bias mainly comes from the characteristics of the sample itself and is not related to the optimization method. Based on the Corollary 1, we can further analyze the performance differences with the guidance of the different types of soft labels. Here we provide Theorem 1 about the relationship between class-wise smoothness degree of soft labels and fairness.
Theorem 1.
Following the setting in Corollary 1, for a dataset $\mathcal{D}$ containing $2$ classes ($c_{+}$ and $c_{-}$), two soft label distribution ($P_{\lambda 1}=\{p_{c_{+}}^{\lambda 1},p_{c_{-}}^{\lambda 1}\}$ and $P_{\lambda 2}=\{p_{c_{+}}^{\lambda 2},p_{c_{-}}^{\lambda 2}\}$) exist, where $P_{\lambda 2}$ have a correct prediction distribution but have a limited different class-wise smoothness degree of soft labels ($v_{1}>0$, $v_{2}>0$):
@@
$p_{c_{-}}^{\lambda 1}(x_{c_{-}})>p_{c_{-}}^{\lambda 2}(x_{c_{-}})=p_{c_{-}}^{\lambda 1}(x_{c_{-}})-v_{2}>0.5,$ (6)
@@
$R(f(x_{c_{+}};\theta_{\lambda 1}))-R(f(x_{c_{-}};\theta_{\lambda 1}))>R(f(x_{c_{+}};\theta_{\lambda 2}))-R(f(x_{c_{-}};\theta_{\lambda 2})).$ (7)
The proof of Theorem 1 can be found in Appendix A.1. Based on Theorem 1, the class-wise smoothness degree of soft labels theoretically has an impact on class-wise robust fairness. If the soft label distribution with different smoothness degree $P_{\lambda 2}$ is applied to guide the model training, where the sharper smoothness degree of soft labels for hard classes and smoother smoothness degree of soft labels for easy classes, the model will appear smaller error risk gap between easy and hard class compared with the soft label distribution with same smoothness degree $P_{\lambda 1}$, which demonstrates better robust fairness. The Theorem 1 theoretically demonstrates that if we appropriately adjust the class-wise smoothness degree of soft labels, the model can achieve class-wise robust fairness.